OpenStudy (anonymous):

Help please @texaschic101 @Abhisar @abb0t @hartnn @IMStuck @sosotiff @quickstudent @kropot72 @myininaya @MrNood @Preetha @amistre64

hartnn (hartnn):

What do you need help with ?

OpenStudy (anonymous):

Given the function \[f (x) =\log_{4}(x + 8), \] find the value of \[f^{-1} (2)\]

OpenStudy (anonymous):

\[\log_4(x+8)=2\] sovle for \(x\) in two steps

hartnn (hartnn):

\(\Large \log_ab=c \implies b=a^c \)

OpenStudy (anonymous):

write in equivalent exponential form as \[4^2=x+8\] and solve for \(x\) (now in one step)

OpenStudy (anonymous):

x = 8

hartnn (hartnn):

thats correct :)

hartnn (hartnn):

oh and thats your \(f^{-1}(2) \\ = 8\)